OpenStudy (josh0404):

can you please help me to solve this problem. if C is the cost, in pesos, of producing x units of the commodity then, assuming that every unit produced is sold, the profit P is given by P=R-C, revenue minus cost, if the cost of producing x units of commodity in (a) is C=975+60x, what is the profit when 60 units are produced (and sold)?

OpenStudy (josh0404):

please help me?

OpenStudy (loser66):

What is R? in other words, R = ??

OpenStudy (josh0404):

Revenue

OpenStudy (loser66):

LOL, of course, it is revenue, but value of it?

OpenStudy (josh0404):

look at the picture Sir.

OpenStudy (loser66):

Common!! That is part c) , I am pretty sure that you have a problem with part a) and b) before c)

OpenStudy (josh0404):

yes sir?

OpenStudy (loser66):

post it, please. We need the value of R.

OpenStudy (josh0404):

ok sir

OpenStudy (josh0404):

Sir did you see it?

OpenStudy (loser66):

Too blur.

OpenStudy (loser66):

But, if you have R from part a), for part c), you just plug x =60 in C , then take P = R -C. Done

OpenStudy (loser66):

\(P = R- (975+60*\color{red}{x})=R-(975+60*\color{red}{60})\)

OpenStudy (josh0404):

Sir wait i ask my classmate to send the clear one

OpenStudy (josh0404):

help me please?